OpenStudy (anonymous):

Hi! I'm confused on how to isolate the y variable. I also need help on writing the standard form of an equation, can anyone help me? Example problem for isolating the y: 2x-3y=4

jhonyy9 (jhonyy9):

2x -3y = 4 so in the first step assum to both sides 3y ,and will get 2x-3y+3y = 4+3y 2x = 4+3y 2. step : subtract from both sides 4 2x-4=4-4+3y 2x-4=3y 3. step : divide both sides by 3 (2x -4)/3 = y hope so much that is understandably good luck